the sum of 1000 is distributed among four people so that each person after the first receives 20 pesos less that the preceding person. How many pesos will each person receive?
1
Expert's answer
2019-08-25T12:40:13-0400

Let the first person receives "x" pesos, then

the second person receives "x-20" pesos,

the third person receives "x-40" pesos,

the fourth person receives "x-60" pesos.

Together they receive "x+x-20+x-40+x-60=4x-120" pesos.

"4x-120=1000,\ 4x=1120,\ x=1120/4=280"

The first person receives "280" pesos , the second person receives "280-20=260" pesos,

the third person receives "280-40=240" pesos and the fourth person receives "280-60=220" pesos.

Answer: The first person receives 280 pesos , the second person receives 260 pesos,

the third person receives 240 pesos and the fourth person receives 220 pesos.